OpenStudy (anonymous):

Solve for x. -3x - 8 + 6x = 7 A. x = - 7/5 B. X = - 1/3 C. X = 5 D. x = 45 Please help :c

OpenStudy (scarlettred):

-3x + 6x = 7 + 8 3x = 15 x = 5

OpenStudy (camerondoherty):

x=5

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

thank you all